OpenStudy (anonymous):

Give an explective rule for the following geometric sequence: A4=-192 A9=-196608 Attempt at Solution: -192=A1(R)^3 -196608=A1(R)^8 -196416=A1(R)^8 + A1(R)^3 ??? Any help greatly appreciated!! Math exam tomorrow.

OpenStudy (asnaseer):

instead of adding, it would be better to divide A9 by A4 because then A1 cancels out.

OpenStudy (anonymous):

Ah so 1024=(r)^8/(r)^5) Now what?

OpenStudy (asnaseer):

now simplify the right-hand-side

OpenStudy (anonymous):

1024=r^3 r=10.079 The real ratio is 4, I dont think I divided the variables with exponents right.

OpenStudy (asnaseer):

no, it should be:\[\frac{r^8}{r^3}=r^5=1024\]

OpenStudy (anonymous):

Oh i see I mistyped when I divided. Thank you so much, I think ive got the hang of this now.

OpenStudy (asnaseer):

great! I'm glad I was able to help :)